OpenStudy (anonymous):

the square root of 48 + 2 times the square root of 3 how do i simplify all of these into one term?

OpenStudy (anonymous):

do you know how to find square root?

OpenStudy (anonymous):

6.9+2*1.7 thats what the equation looks like simplified.

OpenStudy (anonymous):

my teacher told me i need to simplify them into one term

OpenStudy (anonymous):

\[\sqrt{48} +2 \times \sqrt{3}\]

OpenStudy (anonymous):

okay one sec

OpenStudy (anonymous):

6 square root 3

OpenStudy (anonymous):

go to mathway.com and type in your equation ! its super easy and gives you the answers

OpenStudy (anonymous):

k thx